OpenStudy (bri):

the function f(x) = e^x - x^3 has how many critical points?

OpenStudy (anonymous):

What is critical point? o.o

OpenStudy (anonymous):

\[f \prime(x)=e^x-3x^2\] solve f(x)=0 for all x's.. e^x-3x^2=0 is hard to be solved manually. is this a first course in calculus?

OpenStudy (bri):

i know that is what you have to do, i just dont want to do it.

OpenStudy (nowhereman):

let the computer work for you ;-)

OpenStudy (anonymous):

lol.. are you sure of the problem you gave anyway? is it x^3?

OpenStudy (bri):

i typed it correctly. it is e raised to the x minus x raised to the third.

OpenStudy (anonymous):

since you just want to the answer, I used a software program and found that it has three critical points.

OpenStudy (bri):

yeah, i just figured it out because i got tired of waiting, thanks though

OpenStudy (anonymous):

It's not really that hard.. \[e^x - 3x^2 = 0 \rightarrow e^x = 3x^2 \rightarrow x = ln{3} + 2ln{x}\]

OpenStudy (anonymous):

Then just graph both those functions and find where they intersect.

OpenStudy (bri):

wow, polpak, arent you impressive

OpenStudy (anonymous):

Yes, but not because of this.. ;p

OpenStudy (anonymous):

yeah right I didn't think of graphing method.. but I think it's easier to graph e^x=3x^2

OpenStudy (anonymous):

I like the other graph cause e^x grows so fast it's hard to see where they intersect, y=x and lnx are much closer to the origin by comparison.

OpenStudy (anonymous):

^^ good point